A maker of frozen meals claims that the average caloric content of its meals is 400,...

  1. A maker of frozen meals claims that the average caloric content of its meals is 400, and the standard deviation is 15. A researcher tested 16 meals and found that the average number of calories was 350.

    At α = 0.10, which of the following is true?

    a.

    The critical values for the test are ± 1.65.

    b.

    The critical values for the test are ± 1.96.

    c.

    Do not reject the null hypothesis.

    d.

    There is enough evidence to support the claim.
Answer #1

To Test :-
H0 :- µ = 400
H1 :- µ ≠ 400

Reject null hypothesis if | Z | > Z( α/2 )
Critical value Z(α/2) = Z( 0.1 /2 ) = ± 1.65

a.

The critical values for the test are ± 1.65.